OpenStudy (anonymous):

Write a proportion that could be used to solve for each variable. Then solve. 18 ounces in 6 bottles X ounces in 24 bottles

OpenStudy (anonymous):

\[\frac{18}{6}=\frac{x}{24}\\ x=\frac{24\times 18}{6}\]